OpenStudy (anonymous):

find the inverse of the function . f(x)=6x^3 + 3

OpenStudy (mimi_x3):

\[ y= 6x^3 + 3\] \[x = 6y^3 +3\] \[=> 6y^3 = x-3\] \[ y^3 = \frac{x-3}{6}\]

OpenStudy (mimi_x3):

i think that youre able to do it now..

OpenStudy (anonymous):

Let f(x) = y So: \[y = 6x^3 + 3\] \[6x^3 = y - 3\] Divide by 6 both the sides.. \[x^3 = \frac{y}{6} - \frac{1}{2} \implies x = \sqrt[3]{\frac{y}{6} - \frac{1}{2}} \]

OpenStudy (anonymous):

Just flip x and y now and that will be your \(f^{-1}(x) ..\)

OpenStudy (anonymous):

so the final answer is

OpenStudy (anonymous):

f^-1 (x) =3 sq x-3/6

OpenStudy (anonymous):

? is that right?

OpenStudy (anonymous):

Don't you think that will be x/6 ??

OpenStudy (mimi_x3):

abbey is right

OpenStudy (anonymous):

Rest is fine...

OpenStudy (anonymous):

Yes you are right if you put that in brackets..

OpenStudy (mimi_x3):

\[\large f^{-1}x=\sqrt[3]{\frac{x-3}{6}} \]

OpenStudy (anonymous):

If you mean to say this : \[f^{-1}(x) = \sqrt[3]{\frac{x - 3}{6}}\]

OpenStudy (anonymous):

Well Done...

OpenStudy (anonymous):

thank you so much guys!!!!!! i owe u BIG time

OpenStudy (anonymous):

Welcome dear...